Circular No. 7321

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ge IN NGC 309 Y. Sato and W. D. Li, University of California at Berkeley, on behalf

of the Lick Observatory Supernova Search (LOSS; cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Nov. 27.2 (mag ∼ 17.2) and 28.2 UT (mag ∼ 16.9). The new object is located at α = 0^{h}56^{m}43^{s}.76, δ = −9^{o}54′43′′.0 (equinox 2000.0), which is 16′′.0 east and 6′′.7 north of the nucleus of NGC 309. A KAIT image of the same field on Oct. 31.3 (limiting mag ∼ 19.0) showed nothing at the position of SN 1999ge.

SUPERNOVA 1999gf IN UGC 5515 Sato and Li also report the discovery by the LOSS of another apparent

supernova on an unfiltered image taken on Nov. 27.5 UT (mag ∼ 18.2), with confirmation coming from earlier images taken on Nov. 12.5 and 23.5 (when the new object was at mag 17.5 and 18.0, respectively). SN 1999gf is located at α = 10^{h}13^{m}37^{s}.00, δ = −0^{o}56′09′′.6 (equinox 2000.0), which is 18′′.5 west and 37′′.6 south of the nucleus of UGC 5515. A KAIT image taken on 1998 Nov. 21.5 (limiting mag ∼ 19.0) shows nothing at the position of SN 1999gf. UGC 5515 is a member of Abell cluster 957.

SUPERNOVA 1999em IN NGC 1637 P. Garnavich, R. P. Kirshner, and P. Challis, Harvard-Smithsonian

Center for Astrophysics, on behalf of the Supernova Intensive Study collaboration, write: “Spectra of SN 1999em were obtained with the Hubble Space Telescope Imaging Spectrograph on Nov. 5.8 UT. The data were obtained with three low-resolution gratings (range 115–550 nm; resolution ∼ 300 km/s). The combined spectrum has a peak flux of 3 × 10^{−14} erg cm^{−2} s^{−1} A^{−1} at 297 nm, and then drops off rapidly towards the blue, falling a factor of 6 in flux density by 230 nm. The ultraviolet part of the spectrum is characteristic of the ‘Iron Curtain’ resulting from the blending of large number of iron and nickel lines (Hauschildt et al. 1996, Ap.J. 462, 386). A broad P-Cyg feature due to Mg II 280-nm is seen, although modified by the Fe line blanketing. Narrow interstellar absorption lines due to Mg I at 286 nm, Mg II at 280–281 nm, Fe II at 261, 260, 239, and 235 nm, and C IV at 155 nm are also clearly visible at the velocity of the host galaxy.”

1999 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green